# Aerodynamic and structural airfoil shape optimisation via Transfer Learning-enhanced Deep Reinforcement Learning

### v1

@@ 4 Conclusions
While traditional airfoil shape optimisation typically considers aerodynamic efficiency, the optimisation can produce airfoils whose structural integrity might be compromised; in other words, there is a need to preserve some structural constraints –such as the airfoil’s maximum thickness– within the optimisation process. In this paper, we have developed a technique for airfoil shape optimisation –flexible to accomodate single-objective (aerodynamic) or multi-objective (aerodynamic/structural) shape optimisation– that enhances a deep reinforcement learning (DRL) agent with Transfer Learning (TL). Our results indicate that (i) DRL works both for the aerodynamic and aerodynamic/structural shape optimisation problems, (ii) DRL is notably superior to other gradient-free methods such as Particle Swarm Optimisation (PSO) in terms of computational efficiency (being several orders of magnitude faster) and aerodynamic performance, although PSO seem to enforce thickness constraints almost exactly, while DRL does that only approximately, (iii) and the TL-enhancement reduces further training costs with respect to the TL-free case, while approximately maintaining performance (in terms of aerodynamic efficiency and structural preservation).
@@
In summary, our results suggest that DRL stands as an efficient method for airfoil optimisation, both in terms of increasing aerodynamic efficiency and also in the search of geometries that at the same time have enhanced aerodynamic properties and realistically preserve structural constraints. At the same time, the efficiency of this optimisation technique might enable the incorporation of gold standard -yet expensive- aerodynamic solvers, such as CFD-based optimisation, something that also deserves future investigation. Finally, a more general and exciting open problem is to assess whether the benefits of this combined technique also apply in other problems that encompass aerodynamic and structural shape optimisation, e.g. maximisisation of lift-to-drag ratio or flow, minimisation of mass, pressure drop, or stress concentration, and maintaining strength or stiffness, and their application to e.g. pipe shapes, car frames, or fillet optimisation to cite a few.
@@ Acknowledgments
The authors acknowledge funding from project TIFON (PLEC2023-010251) funded by
MCIN/AEI/10.13039/501100011033, Spain.
The authors acknowledge funding from the European Union (project HERFUSE) under GA No 101140567. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or Clean Aviation Joint Undertaking. Neither the European Union nor Clean Aviation JU can be held responsible for them.
LL acknowledges partial financial support from via project MISLAND
(PID2020-114324GB-C22)
and the Maria de Maeztu Program for units of Excellence (grant CEX2021-001164-M), both funded by MICIU/AEI/10.13039/501100011033, and funding from the European Commission Chips Joint Undertaking project No. 101194363 (NEHIL).
GR acknowledges partial financial support received by the Grant DeepCFD (Project No. PID2022-137899OB-I00) funded by MICIU/AEI/10.13039/501100011033 and by ERDF, EU.
Finally, all authors gratefully acknowledge the Universidad Politécnica de Madrid (www.upm.es) for providing computing resources on Magerit Supercomputer.
@@ Appendix A RL environment configuration
This section shows the configuration used for the environments.
A.1 CST parameters
@@
| Upper surface | 8 x [-1.5] | 8 x [1.25] |
| Lower surface | 8 x [-0.75] | 8 x [1.5] |
| Trailing-edge thickness | 0.0005 | 0.01 |
| Leading-edge weight | -0.05 | 0.775 |
The initial airfoil for an episode is selected from this list of 20 NACA airfoils: naca0006, naca0009, naca0012, naca0015, naca0018, naca1408, naca1410, naca1412, naca2412, naca2415, naca4412, naca4415, naca4420, naca6412, naca6415, naca7421, naca8409, naca8412, naca8415, naca9421
@@ Appendix B PPO hyperparameters
If a parameter is not included in a table, the default value given by Stable Baselines 3 will be used.
B.1 Improvement of training efficiency
Sample inefficiency is a common problem for many reinforcement learning algorithms, which means that to learn, the agent needs to see too many simulations. For problems like this, where the simulations are expensive, this could be an issue. To reduce the number of simulations needed and thus the computational cost for the training process, the hyperparameters of PPO have been tuned following this goal. After many trainings, it has been observed that the most relevant parameter in terms of the number of steps needed to train an agent is the discount factor, , and the second one, the clip range. Moreover, it was found empirically that the optimum value of is 0.3, and the value of clip range was increased, to a greater or lesser extent, from its default value 0.2. For instance, with this hyperparameter tuning and using Neuralfoil, an agent can be trained on 60000 steps that has the same performance as one trained on 120000 steps using the default hyperparameters given by Stable Baselines 3. Figure 7 shows the mean reward that the agent obtained during training for different values of . As can be observed, the line for grows faster than the line for , which means that the agent trained with is learning with fewer steps.
B.2 For agent trained with XFoil
| Parameter | Value |
| Total timesteps | - |
| Learning rate | |
| Steps per update | 2048 |
| Batch size | 64 |
| Num. epochs | 20 |
| Discount factor () | 0.3 |
| Clip range | 0.3 |
| Entropy coefficient | 0.001 |
| Policy net architecture | [256, 256] |
| Value net architecture | [256, 256] |
B.3 Fine-tuning Neuralfoil and XFoil
| Parameter | Value |
| Total timesteps | 26312 |
| Learning rate | |
| Steps per update | 2048 |
| Batch size | 64 |
| Num. epochs | 10 |
| Discount factor () | 0.3 |
| Clip range | 0.6 |
| Entropy coefficient | 0 |
| Policy net architecture | [256, 256] |
| Value net architecture | [256, 256] |
| Parameter | Value |
| Total timesteps | 10240 |
| Learning rate | |
| Steps per update | 512 |
| Batch size | 64 |
| Num. epochs | 20 |
| Discount factor () | 0.3 |
| Clip range | 0.2 |
| Entropy coefficient | 0.005 |
| Policy net architecture | [256, 256] |
| Value net architecture | [256, 256] |

### v2

@@ IV Conclusions
While traditional airfoil shape optimisation typically considers aerodynamic efficiency, the optimisation can produce airfoils whose structural integrity might be compromised; in other words, there is a need to preserve some structural constraints –such as the airfoil’s maximum thickness– within the optimisation process. In this paper, we have developed a technique for airfoil shape optimisation –flexible to accomodate purely aerodynamic or hybrid (aerodynamic/structural) shape optimisation– that enhances a deep reinforcement learning (DRL) agent with Transfer Learning (TL). Our results indicate that (i) DRL works both for the aerodynamic and aerodynamic/structural shape optimisation problems, (ii) DRL is notably superior to other gradient-free methods such as Particle Swarm Optimisation (PSO) in terms of computational efficiency (being several orders of magnitude faster) and aerodynamic performance, although PSO seem to enforce thickness constraints almost exactly, while DRL does that only approximately, (iii) and the TL-enhancement reduces further training costs with respect to the TL-free case around a 86%, while approximately maintaining performance (in terms of aerodynamic efficiency and structural preservation).
In summary, our results suggest that DRL stands as an efficient method for airfoil optimisation, both in terms of increasing aerodynamic efficiency and also in the search of geometries that at the same time have enhanced aerodynamic properties and realistically preserve structural constraints.
@@
On relation to the use of TL-enhanced DRL, our results encourage the exploration of more challenging problems, such as e.g. the optimisation of (3-dimensional) wings: a problem where both the computational cost of the solvers and the complexity of the shape parameterisation increase significantly, and where transfer learning techniques that use inexpensive surrogate solvers could be of great help. More generally, the efficiency of this optimisation technique might enable the incorporation of gold standard -yet expensive- aerodynamic solvers, such as CFD-based optimisation, something that also deserves future investigation. Finally, a more general and exciting open problem is to assess whether the benefits of this combined technique also apply in other problems that encompass aerodynamic and structural shape optimisation, e.g. maximisisation of lift-to-drag ratio or flow, minimisation of mass, pressure drop, or stress concentration, and maintaining strength or stiffness, and their application to e.g. pipe shapes, car frames, or fillet optimisation to cite a few. Other interesting applications where DRL could be of great use include the design of propellers, rotor blades, and turbine components in aerospace; aerodynamic shaping of vehicles and structural frames in automotive engineering; and flow optimisation in energy systems like HVAC ducts, wind farms, and hydraulic networks.
On relation to the aerodynamic vs structural optimisation scenario, in this work we initially chose to scalarize the initial multi-objective optimisation problem by combining the aerodynamic and the structural integrity objectives into a single reward function. For illustration, we advance in Fig. 7 the results of a Pareto front analysis, where the values of both aerodynamic and structural integrity improvements are plotted for a variety of values of the hyperparameter . Further work is needed to consider both aerodynamic and structural terms independently as well as to add more sophisticated structural criteria, in order to embrace full-fledged multi-objective optimisation.
Supplementary Material
Supplementary material contains URLs where animations of the optimisation process of the airfoils in table 1 are available for the different algorithms described.
@@ Acknowledgments
The authors thank useful discussions within the NUMATH group. DR, EV and GR acknowledge funding from project TIFON (PLEC2023-010251) funded by MCIN/AEI/10.13039/501100011033, Spain, funding from the European Union (project HERFUSE) under GA No 101140567. Views and opinions expressed are those of the authors only, and do not necessarily reflect those of the European Union or Clean Aviation Joint Undertaking. Neither the European Union nor Clean Aviation JU can be held responsible for them. LL acknowledges partial financial support from via project MISLAND (PID2020-114324GB-C22), CSxAI (PID2024-157526NB-I00) and the Maria de Maeztu Program for units of Excellence (Grant No. CEX2021-001164 M), the three of them funded by MICIU/AEI/10.13039/501100011033, and funding from the European Commission Chips Joint Undertaking project No. 101194363 (NEHIL). GR acknowledges partial financial support received by the Grant DeepCFD (Project No. PID2022-137899OB-I00) funded by MICIU/AEI/10.13039/501100011033 and by ERDF, EU. Finally, all authors gratefully acknowledge the Universidad Politécnica de Madrid (www.upm.es) for providing computing resources on Magerit Supercomputer.
@@ Appendix A RL environment configuration
This section shows the configuration used for the environments. The ranges of the CST parameters, shown in Table 7, were selected based on empirical testing, aiming to strike a balance between allowing the agent sufficient freedom to explore the design space and avoiding excessively extreme or unrealistic geometries.
@@
| Upper surface | ||
| Lower surface | ||
| Trailing-edge thickness | ||
| Leading-edge weight |
The initial airfoil for an episode is selected from this list of 20 NACA airfoils: naca0006, naca0009, naca0012, naca0015, naca0018, naca1408, naca1410, naca1412, naca2412, naca2415, naca4412, naca4415, naca4420, naca6412, naca6415, naca7421, naca8409, naca8412, naca8415, naca9421.
@@ Appendix B PPO hyperparameters
This section provides a summary of the PPO hyperparameters used to train the different agents discussed in this work. Sample inefficiency is a common problem for many reinforcement learning algorithms, which means that to learn, the agent needs to see too many simulations. For problems like this, where the simulations are expensive, this could be an issue. To minimise the number of required simulations and consequently lower the computational cost of the training process, the PPO hyperparameters have been adjusted with this objective in mind. After many trainings where these hyperparameters were being modified, it has been observed, as expected, that the most relevant parameter in terms of the number of steps needed to train an agent is the discount factor, , and the second one, the clip range. Moreover, it was found empirically that the optimum value of is 0.3, and the value of clip range was increased, to a greater or lesser extent, from its default value 0.2. For instance, with this hyperparameter tuning and using Neuralfoil, an agent can be trained on 60000 steps that has the same performance as one trained on 120000 steps using the default hyperparameters given by Stable Baselines 3. Since our primary goal was to reduce training cost through transfer learning, we did not perform an exhaustive hyperparameter search and considered the empirical tuning sufficiently effective for our purposes. Nonetheless, we conducted a hyperparameter search using Optuna Akiba et al. (2019) for the NeuralFoil agent. It is worth noting that the resulting configuration closely aligns with the parameters previously obtained through empirical tuning. Figure 8 shows the mean reward that the agent obtained during training for different values of . As can be observed, the line for grows faster than the line for , which means that the agent trained with is learning with fewer steps.
Table 8 below consolidates the PPO hyperparameters for three training configurations: one agent trained from scratch using XFoil, the pre-trained agent with NeuralFoil and the fine-tuned from a previously trained XFoil policy. If a parameter is not included in the table, the default value given by Stable Baselines 3 will be used.
| Solver | XFoil | NeuralFoil | XFoil |
| TL strategy | from scratch | pre-training | transfer learning |
| Total timesteps | - | 26312 | 10240 |
| Learning rate | |||
| Steps per update | 2048 | 2048 | 512 |
| Batch size | 64 | 64 | 64 |
| Num. epochs | 20 | 10 | 20 |
| Discount factor () | 0.3 | 0.3 | 0.3 |
| Clip range | 0.3 | 0.6 | 0.2 |
| Entropy coefficient | 0.001 | 0 | 0.005 |
| Policy net arch. | [256, 256] | [256, 256] | [256, 256] |
| Value net arch. | [256, 256] | [256, 256] | [256, 256] |
@@ Appendix C XFoil configuration
This section details the configuration settings used for XFoil during the generation of simulation data for training and evaluation. The selected parameters, shown in Table 9, were chosen to ensure a balance between computational efficiency and numerical stability.